# Auriga Streams I: disrupting satellites surrounding Milky Way-mass haloes at multiple resolutions

### v1

@@ 3.2 Classifying systems
We only consider accretion events with at least 100 star particles to ensure that the distribution of each object is adequately captured. This limit is more conservative than what is typically used to study satellite systems in cosmological simulations ($\gtrsim$10 bound star particles; Sawala et al., 2016; Simpson et al., 2018), since we need to study the phase-space distribution of accretions as they disrupt. Given the typical baryonic mass resolution at each simulation level, in addition to stellar winds that reduce the mass of star particles over time, this limit corresponds to a minimum total stellar mass of 3.7×$\times 10^{6}$ M, 4.8×$\times 10^{5}$ M, and 5.7×$\times 10^{4}$ M at levels 4, 3, and 2 respectively. This also means that we effectively ignore accreted material from objects below this limit, which typically correspond to $\lesssim$0.5 per cent of accreted stars, but can be as high as 3 per cent for a handful of level 4 haloes.
@@ 3.2.1 Bound fraction
Figure 2: Top: fraction of stellar mass still bound to the progenitor, $f_{\text{bound}}$, for all accretions in the visually classified sample, split according to visual classification. Inset: zoom on higher $f_{\text{bound}}$ for the subset of objects that were classified by eye as intact satellites. We use $f_{\text{bound}}=0.97$ (dashed vertical line) as the boundary between intact satellite and stream for our automated classification. Bottom: distribution of $f_{\text{bound}}$ for all accretions, split by classification assigned automatically. Shaded regions indicate objects that have $10<f_{\text{bound}}<$ 1, which combined with the right-most bin of $f_{\text{bound}}=1$ systems would traditionally be treated as ‘satellites’ in simulation studies. The majority of satellites with a bound progenitor have lost substantial amounts of stellar mass.
The distinction between intact satellites and stellar streams is based on how much stellar mass the progenitor has lost by the present day. The fraction of the object’s stellar mass that is bound to the progenitor is determined by SUBFIND and only includes stellar mass formed in that object (i.e. excluding other mergers whose debris is now bound to that SUBFIND subhalo). The top panel of Figure 2 shows the distribution of $f_{\text{bound}}$ for our visually classified sample. Nearly all intact satellites have $f_{\text{bound}}\sim 1$, while the few instances that are substantially below this are typically low-mass systems that have only lost a few star particles (out of 100-200 total). We set the threshold at $f_{\text{bound}}=0.97$, which captures nearly all structures that were visually classified as intact (inset panel of Figure 2). In addition, when sorting the full sample by descending $f_{\text{bound}}$, the transition from where most systems were classified as intact (with a few stream interlopers) to where most systems were classified as streams (with a few intact interlopers) also occurs near $f_{\text{bound}}=0.97$.
@@ 3.2.2 Local velocity dispersion
which scales the particles’ velocities with a weighting factor $w_{\text{v}}$ that has units of kpc km-1 s, such that $\Delta$ has units of kpc. The local velocity dispersion is then the velocity dispersion among the $k$ neighbours with the smallest $\Delta$, where $k$ is taken to be 1 per cent of the total number of star particles in the object or a minimum of 7 particles, whichever is larger. We repeat this exercise to compute the local velocity dispersion for all star particles in the same accretion event, then take the median of this distribution for our classification purposes (see below). Fixing the fraction of particles in the accretion, rather than the number of neighbours $k$, controls for the increased number of particles (and therefore smaller amount of phase space occupied by each particle) for the same structure at higher resolution levels. The absolute minimum of 7 particles prevents a biased calculation of the local velocity dispersion based on too few neighbours and is the same absolute minimum adopted in Panithanpaisal et al. (2021).
@@
$\sigma^{50}_{\text{vel}}=A\log\left(\frac{M_{\star}}{M_{\odot}}\right)+B,$ (2)
where $\sigma^{50}_{\text{vel}}$ is the median of the local velocity dispersion and $M_{\star}$ is the total stellar mass. We run this SVM step on the full visually classified set of streams and phase-mixed systems, including data at all three resolution levels. In our automatic classification of all systems, accretions with $f_{\text{bound}}<0.97$ that fall above this line are classified as phase-mixed, while those below the line are classified as streams. Intact satellites nearly always have the lowest $\sigma^{50}_{\text{vel}}$ for their stellar mass. There is a modest positive slope on the mass dependence of Equation 2, as well as a minimum possible median local velocity dispersion that grows with stellar mass, which reflects the fact that more massive accretions have higher velocity dispersions prior to infall.
@@
Finally, we caution that our exact classification algorithm is not intended to be plug-and-play for analysis of other simulation suites. For example, the $f_{\text{bound}}=0.97$ intact/stream separation determined in Section 3.2.1 could have arisen from particular minutae of the Auriga model or SUBFIND halo finder. However, the general approach should be broadly applicable, and the modifications we have made to the method outlined in Panithanpaisal et al. (2021) are likely to be useful for any analysis that considers simulations at different mass resolutions.
@@ 4 Matching across resolutions
To assign a best match for each debris group in the low resolution simulation, we find the pairing with the high resolution simulation that has the smallest value of $\Phi$. If this pairing has a value $\Phi<1$ and the particle ratio is within a factor of 2 of 8 (dashed box in 4) the group is assigned to be a ‘match’. In cases where the minimum $\Phi$ corresponding group lies outside this region, the group is assigned to be a ‘marginal match’. If multiple low-resolution groups are marginally matched to the same high-resolution group, we label the most massive low-resolution object as the marginal match and all other objects are labelled ‘unmatched’.
@@
We considered several matching techniques and initially tried the method described in Grand et al. (2021), however that method was applied to simulations that had the same number of snapshot outputs. Five of our level 3 resolution simulations have only 64 snapshots, instead of the 128 snapshots for levels 4 and 2. We found the method of Grand et al. (2021) did not produce good matches in these cases. We also considered matching dark matter particles to their nearest neighbour in initial conditions and comparing debris groups on a particle-by-particle basis. However, we found that this more complex technique did not give better matches than our adopted approach.
Across the six halos with both level 4 and level 3 resolution runs, 90 per cent of intact satellites, over 98 per cent of streams, and over 91 per cent of phase-mixed systems present in level 4 were matched by our method. The fraction of systems in the three categories that were not matched but are present in the level 4 runs were 4, 6, and 1 per cent respectively, for a total of 8 objects across all 6 halos. There were systems that were present in the level 3 simulations that were not matched to a level 4 systems, but this is expected due to resolution effects of the simulations (Grand et al., 2021). For the level 2 run of Au-6, all intact satellites and streams that were present in the level 3 simulation were present in the level 2 simulation. Only 1 phase-mixed system in the level 3 simulation was not matched in the level 2 simulation.
@@
Figure 6: Stellar morphologies of individual structures in Au-6 matched across three resolution levels (increasing in resolution from top to bottom). The colour corresponds to the amount of stellar mass in a pixel, logarithmically scaled to improve contrast in low/high-density regions (the normalisation is consistent across resolution but not for different columns, as the objects span a wide range in stellar mass). Cases where the morphology does not match across all levels (third from the left, third from the right) correspond to the same accretion event experiencing slightly different accretion timing (see Paper II for further detail). The stellar morphologies of accretions in Auriga are remarkably converged across resolution, often localizing the progenitor to a similar phase along its orbit.
@@ 5 The three-phase structure of accreted stellar mass
Figure 8: Left: number of accretion events in each halo, split by classification. Right: fraction of accreted stellar mass in each classification. Haloes are ordered within a given level according to the fraction of accreted stellar mass that is in a phase-mixed system; ordering is consistent between left and right panels. Information for specific haloes is listed in Table 1. Vertical dashed lines separate resolution levels, with 28, 6, and 1 halo(es) in levels 4, 3, and 2, respectively. While the number of systems in each category is roughly consistent across hosts, the fraction of stellar mass contributed by each phase of disruption spans a wide range of possibilities.
@@
In Figure 9 we also recover the familiar result that most of the accreted stellar mass comes from one (or a few) massive accretion(s). This is a natural outcome of the halo mass function set by $\Lambda$CDM and the steepness of the stellar mass-halo mass relation that has been validated by both semi-analytic modeling (Bullock & Johnston, 2005; Cooper et al., 2010; Deason et al., 2016; Amorisco, 2017) and hydrodynamic simulations (Monachesi et al., 2019; Santistevan et al., 2020; Fattahi et al., 2020; Wright et al., 2024). Perhaps unsurprisingly, the right panel of Figure 9 demonstrates that this picture holds across the three morphological types, with the overall contribution from intact, stream, and phase-mixed morphologies each dominated by a few massive objects.
@@ 6.1 Most satellites are streams
One of the most striking results from our classification analysis is how prevalent streams are in our catalogue of accreted material. Streams outnumber intact satellites by a factor of two (Figure 9) and most objects with a bound progenitor, which would traditionally be labelled as ‘satellites’, have experienced substantial amounts of disruption (Section 3.2.1 and Figure 2). Depending on the exact threshold for $f_{\text{bound}}$, approximately 50 to 70 per cent of progenitors in Auriga are in the process of actively disrupting. As a sort of inversion of this result, 80 per cent of Auriga streams still have a progenitor ($f_{\text{bound}}>0$) at the present day (74 per cent for level 4, 91 per cent for level 3).
Cosmological simulations based on $\Lambda$CDM using both semi-analytic (Li et al., 2010; Barber et al., 2015) and hydrodynamic (Brooks & Zolotov, 2014; Wetzel et al., 2016; Sawala et al., 2016; Fattahi et al., 2018) prescriptions consistently invoke tidal disruption to explain properties of the observed Milky Way satellites such as the $M_{\ast}-V_{\text{max}}$ relation. Specifics on the amount of stellar mass lost on an object-by-object basis are rarer. Figure 12 of Cooper et al. (2010) shows the distribution of $1-f_{\text{bound}}$ for six Aquarius haloes, which bears a striking resemblance to the bottom panel of our Figure 2. Using the APOSTLE simulations, Wang et al. (2017) found that $\sim$30 per cent of satellites lose more than 20 per cent of their stellar mass after infall, though the central stellar discs in Auriga are more massive than in APOSTLE, affecting the number (Richings et al., 2020) and orbits (Riley et al., 2019) of bound satellites. Shipp et al. (2023) found that 61 out of 140 satellites with bound mass $M_{\ast}\gtrsim 5\times 10^{5}$ M, or 44 per cent, are progenitors of streams identified in FIRE by Panithanpaisal et al. (2021).
These results are striking, considering that of the 13 Milky Way satellites that are above $M_{\ast}=5\times 10^{5}$ M,  [^9] Based on the Local Volume Database. only Sagittarius shows clear tidal tails, along with Antlia II that has a large size and velocity gradient indicative of tidal disruption (Torrealba et al., 2019; Ji et al., 2021). In addition, there are only two known Milky Way streams (Sagittarius and OC; Li et al., 2022) originating from satellite galaxies above the same mass scale. However, most of the streams predicted by cosmological simulations are very low surface brightness and would be detected as bound progenitors or evade detection altogether by current imaging surveys, but may be revealed by the upcoming facilities (Wang et al., 2017; Shipp et al., 2023). Ultimately, assessing the consistency of the Auriga streams with current observations and making predictions for future surveys will require detailed mock observations in the style of Shipp et al. (2023) for the Milky Way and Miro-Carretero et al. (2024b) for extragalactic hosts.
@@ 6.2 Streams formed from preprocessing
Figure 11: Example of a stream that is disrupting around another satellite as it accretes onto the main host. The background colour represents the amount of stellar mass from the primary satellite in a pixel, logarithmically scaled. Individual points correspond to star particles from the disrupting secondary satellite. The dashed line traces the recent trajectory of the primary satellite and the cross indicates the centre of the main host (note that the centre of the primary is 310 kpc from the host and apparent proximity of the tidal debris to the cross is a projection effect). Systems merging with each other prior to falling into the main host, and the streams that naturally form as a result, are included in our sample.
@@
These preprocessed systems can appear as outliers in several parameters; objects classified as streams or phase-mixed with recent ($<2$ Gyr) accretion times onto the main host and large pericentres (see Paper II for more detail). While we reserve a more in-depth look at stellar streams formed via group infall for future work, it can be useful to have a simple flag to indicate potential preprocessing. We trace every system backwards in time to see if there are three consecutive snapshots where it is not the most massive halo in its FOF group, ignoring cases where the FOF group is that of the main Milky Way-mass host. If this condition is satisfied, we label that system as a case of preprocessing [^10] While we do not examine whether an accretion event actually begins disrupting before accreting onto the main host, this definition does capture all of the instances of preprocessing that we identified visually or as clear outliers in accretion time or dynamical quantities (see Paper II).. Across all haloes and resolution levels, 17 per cent of accretion events satisfy this condition, with an increased prevalence of 28 per cent for objects classified as streams (18 per cent for intact, 11 per cent for phase-mixed). These fractions agree with those computed for levels 4 and 3 separately, except the value for intact satellites is only 4 per cent in level 3, possibly reflecting the unique satellite interactions for which those haloes were selected.
Clearly preprocessing of satellites plays a substantial role in producing stellar streams, an effect which has been relatively under-explored. We emphasise that this mechanism is not rare, with approximately 1 in 4 streams in our sample residing in a more massive halo prior to infall. It shines light on a tantalising question – if the Milky Way ($M_{\text{200c}}\sim 1\times 10^{12}$ M, $M_{\ast}\sim 5\times 10^{10}$ M; Bland-Hawthorn & Gerhard, 2016) has over 100 stream candidates with 8 streams spectroscopically confirmed to originate from satellite galaxies, where are the streams disrupting around the LMC ($M_{\text{200c}}\sim 2\times 10^{11}$ M, $M_{\ast}\sim 3\times 10^{9}$ M; Vasiliev, 2023)? Future deep photometry from the Rubin Observatory and Roman Space Telescope could provide exciting answers in the coming decade, potentially adding a new layer to the tangled web of disrupting satellites that comprise the Galactic halo.

### v2

@@ 3.2 Classifying systems
[…]
@@ 3.2.1 Bound fraction
Figure 2: Top: fraction of stellar mass still bound to the progenitor, $f_{\text{bound}}$, for all accretions in the visually classified sample, split according to visual classification. Inset: zoom on higher $f_{\text{bound}}$ for the subset of objects that were classified by eye as intact satellites or streams. We use $f_{\text{bound}}=0.97$ (dashed vertical line) as the boundary between intact satellite and stream for our automated classification. Bottom: distribution of $f_{\text{bound}}$ for all accretions, split by classification assigned automatically. Shaded regions indicate objects that have $0<f_{\text{bound}}<1$, which combined with the right-most bin of $f_{\text{bound}}=1$ systems would traditionally be treated as ‘satellites’ in simulation studies. The majority of satellites with a bound progenitor have lost substantial amounts of stellar mass.
The distinction between intact satellites and stellar streams is based on how much stellar mass the progenitor has lost by the present day. The fraction of the object’s stellar mass that is bound to the progenitor is determined by SUBFIND and only includes stellar mass formed in that object (i.e. excluding other mergers whose debris is now bound to that SUBFIND subhalo). The top panel of Figure 2 shows the distribution of $f_{\text{bound}}$ for our visually classified sample. Nearly all intact satellites have $f_{\text{bound}}\sim 1$ (across all haloes and resolution levels, 58 objects have $f_{\text{bound}}=1$ and 84 have $0.99<f_{\text{bound}}<1$), while the few instances that are substantially below this are typically low-mass systems that have only lost a few star particles (out of 100-200 total). We set the threshold at $f_{\text{bound}}=0.97$, which captures nearly all structures that were visually classified as intact (inset panel of Figure 2). In addition, when sorting the full sample by descending $f_{\text{bound}}$, the transition from where most systems were classified as intact (with a few stream interlopers) to where most systems were classified as streams (with a few intact interlopers) also occurs near $f_{\text{bound}}=0.97$.
@@ 3.2.2 Local velocity dispersion
which scales the particles’ velocities with a weighting factor $w_{\text{v}}$ that has units of kpc km-1 s, such that $\Delta$ has units of kpc. The local velocity dispersion is then the velocity dispersion among the $k$ neighbours with the smallest $\Delta$, where $k$ is taken to be 1 per cent of the total number of star particles in the object or a minimum of 7 particles, whichever is larger. We repeat this exercise to compute the local velocity dispersion for all star particles in the same accretion event, then take the median of this distribution for our classification purposes (see below). Fixing the fraction of particles in the accretion, rather than the number of neighbours $k$, controls for the increased number of particles (and therefore smaller amount of phase space occupied by each particle) for the same structure at higher resolution levels. The absolute minimum of 7 particles prevents a biased calculation of the local velocity dispersion based on too few neighbours and is the same absolute minimum adopted in Panithanpaisal et al. (2021). The adopted value of 1 per cent is relatively arbitrary – larger values push the definition of ‘local’ further away in phase space from the particle of interest, while smaller values result in more objects falling back to the minimum of 7. Across the full sample, $\sim$55 per cent of objects are under the 1 per cent regime (i.e. have more than 700 star particles).
@@
$\sigma^{50}_{\text{vel}}=A\log\left(\frac{M_{\ast}}{M_{\odot}}\right)+B,$ (2)
where $\sigma^{50}_{\text{vel}}$ is the median of the local velocity dispersion and $M_{\ast}$ is the total stellar mass. We run this SVM step on the full visually classified set of streams and phase-mixed systems, including data at all three resolution levels. In our automatic classification of all systems, accretions with $f_{\text{bound}}<0.97$ that fall above this line are classified as phase-mixed, while those below the line are classified as streams. Intact satellites nearly always have the lowest $\sigma^{50}_{\text{vel}}$ for their stellar mass. There is a modest positive slope on the mass dependence of Equation 2, as well as a minimum possible median local velocity dispersion that grows with stellar mass, which reflects the fact that more massive accretions have higher velocity dispersions prior to infall.
@@
Finally, we caution that our exact classification algorithm (the values of $w_{v}$, $A$, $B$, $k$ listed above) is not intended to be plug-and-play for analysis of other simulation suites. For example, the $f_{\text{bound}}=0.97$ intact/stream separation determined in Section 3.2.1 could have arisen from particular minutae of the Auriga model or SUBFIND halo finder. However, the general approach (a classification based on boundaries in $f_{\text{bound}}$ and $M_{\ast}$ vs. $\sigma_{\text{vel}}^{50}$) should be broadly applicable, and the modifications we have made to the method outlined in Panithanpaisal et al. (2021) are likely to be useful for any analysis that considers simulations at different mass resolutions.
@@ 4 Matching across resolutions
To assign a best match for each debris group in the low resolution simulation, we find the pairing with the high resolution simulation that has the smallest value of $\Phi$. If this pairing has a value $\Phi<1$ and the particle ratio is within a factor of 2 of 8 (dashed box in 4) the group is assigned to be a ‘match’. In cases where the minimum $\Phi$ corresponding group lies outside this region, the group is assigned to be a ‘marginal match’. If multiple low-resolution groups are marginally matched to the same high-resolution group, we label the most massive low-resolution object as the marginal match and all other objects are labelled ‘unmatched’. There are no cases where multiple low-resolution groups are high-quality matches to the same high-resolution group.
@@
We considered several matching techniques and initially tried the method described in Grand et al. (2021), which finds the high resolution counterpart that minimizes the sum of positional offsets across all snapshots prior to infall. However, that method was applied to simulations that had the same number of snapshot outputs. Five of our level 3 resolution simulations have only 64 snapshots, instead of the 128 snapshots for levels 4 and 2. We found the method of Grand et al. (2021) did not produce good matches in these cases. We also considered matching dark matter particles to their nearest neighbour in initial conditions and comparing debris groups on a particle-by-particle basis. However, we found that this more complex technique did not give better matches (as determined by visual comparison of dark matter distributions at $z=0$ and host-centric distance vs. time) than our adopted approach.
Across the six halos with both level 4 and level 3 resolution runs, 27 intact satellites, 67 streams, and 108 phase-mixed systems present in level 4 were matched by our method. The number of systems in the three categories that were not matched (either ‘marginal’ or ‘unmatched’) the level 4 runs were 4, 1, and 10 respectively, a match rate of over 90 per cent for each morphology across all 6 halos. There were systems present in the level 3 simulations that were not matched to a level 4 systems, but this is expected due to resolution effects of the simulations (Grand et al., 2021). For the level 2 run of Au-6, all intact satellites and streams that were present in the level 3 simulation were present in the level 2 simulation. Only 1 phase-mixed system in the level 3 simulation was not matched in the level 2 simulation.
@@
Figure 6: Stellar morphologies of individual structures in Au-6 matched across three resolution levels (increasing in resolution from top to bottom). We provide the classification label from Section 3.2; objects labelled ‘unclassified’ have fewer than 100 star particles. The shade corresponds to the amount of stellar mass in a pixel, logarithmically scaled to improve contrast in low/high-density regions (the normalisation is consistent across resolution but not for different columns, as the objects span a wide range in stellar mass). Cases where the morphology does not match across all levels (third from the left, third from the right) correspond to the same accretion event experiencing slightly different accretion timing (see Paper II for further detail). The stellar morphologies of accretions in Auriga are remarkably converged across resolution, often localizing the progenitor to a similar phase along its orbit.
@@ 5 The three-phase structure of accreted stellar mass
Figure 8: Left: number of accretion events in each halo, split by classification. Right: fraction of accreted stellar mass in each classification. Haloes are ordered within a given level according to the fraction of accreted stellar mass that is in a phase-mixed system; ordering is consistent between left and right panels. Information for specific haloes is listed in Table 1. Vertical dashed lines separate resolution levels, with 28, 6, and 1 halo(es) in levels 4, 3, and 2, respectively. Haloes that are re-simulated at higher resolution are specified to assist comparison. While the number of systems in each category is roughly consistent across hosts, the fraction of stellar mass contributed by each phase of disruption spans a wide range of possibilities.
@@
[…]
@@ 6.1 Most satellites are streams
One of the most striking results from our classification analysis is how prevalent streams are in our catalogue of accreted material. Streams outnumber intact satellites by a factor of two (Figure 9) and most objects with a bound progenitor, which would traditionally be labelled as ‘satellites’, have experienced substantial amounts of disruption (Section 3.2.1 and Figure 2). Depending on the exact threshold for $f_{\text{bound}}$, approximately 50 to 70 per cent of progenitors in Auriga are in the process of actively disrupting. As a sort of inversion of this result, 80 per cent of Auriga streams still have a progenitor ($f_{\text{bound}}>0$) at the present day (74 per cent for level 4, 91 per cent for level 3, see Section 6.5 for discussion on numerical convergence in $f_{\text{bound}}$ across levels).
[…]
These results are striking, considering that of the 13 Milky Way satellites that are above $M_{\ast}=5\times 10^{5}$ M,  [^9] Based on the Local Volume Database (Pace, 2024). only Sagittarius shows clear tidal tails, along with Antlia II that has a large size and velocity gradient indicative of tidal disruption (Torrealba et al., 2019; Ji et al., 2021). In addition, there are only two known Milky Way streams (Sagittarius and OC; Li et al., 2022) originating from satellite galaxies above the same mass scale [^10] To arrive at this tally, we adopt metallicity measurements from Li et al. (2022), Chenab’s metallicity for the OC stream, and a cutoff at [Fe/H] =1.8=-1.8 (Kirby et al., 2013). Relaxing any of these criteria could reasonably lead to the inclusion of the Jhelum, Cetus-Palca, and Indus streams (e.g. for higher mass estimates for Cetus-Palca, see Thomas & Battaglia, 2022; Yuan et al., 2022).. However, most of the streams predicted by cosmological simulations are very low surface brightness and would be detected as bound progenitors or evade detection altogether by current imaging surveys, but may be revealed by the upcoming facilities (Wang et al., 2017; Shipp et al., 2023). Ultimately, assessing the consistency of the Auriga streams with current observations and making predictions for future surveys will require detailed mock observations in the style of Shipp et al. (2023) for the Milky Way and Miró-Carretero et al. (2025) for extragalactic hosts.
@@ 6.2 Streams formed from preprocessing
Figure 11: Example of a stream that is disrupting around another satellite as it accretes onto the main host. The background shade represents the amount of stellar mass from the primary satellite in a pixel, logarithmically scaled. Individual points correspond to star particles from the disrupting secondary satellite. The dashed line traces the recent trajectory of the primary satellite and the cross indicates the centre of the main host (note that the centre of the primary is 310 kpc from the host and apparent proximity of the tidal debris to the cross is a projection effect). Systems merging with each other prior to falling into the main host, and the streams that naturally form as a result, are included in our sample.
@@
These preprocessed systems can appear as outliers in several parameters; objects classified as streams or phase-mixed with recent (<2<2 Gyr) accretion times onto the main host and large pericentres (see Paper II for more detail). While we reserve a more in-depth look at stellar streams formed via group infall for future work, it can be useful to have a simple flag to indicate potential preprocessing. We trace every system backwards in time to see if there are three consecutive snapshots where it is not the most massive halo in its FOF group, ignoring cases where the FOF group is that of the main Milky Way-mass host. If this condition is satisfied, we label that system as a case of preprocessing [^11] While we do not examine whether an accretion event actually begins disrupting before accreting onto the main host, this definition does capture all of the instances of preprocessing that we identified visually or as clear outliers in accretion time or dynamical quantities (see Paper II).. Across all haloes and resolution levels, 17 per cent of accretion events satisfy this condition, with an increased prevalence of 28 per cent for objects classified as streams (18 per cent for intact, 11 per cent for phase-mixed). These fractions agree with those computed for levels 4 and 3 separately, except the value for intact satellites is only 4 per cent in level 3, possibly reflecting the unique satellite interactions for which those haloes were selected.
Clearly preprocessing of satellites plays a substantial role in producing stellar streams, an effect which has been relatively under-explored. We emphasise that this mechanism is not rare, with approximately 1 in 4 streams in our sample residing in a more massive halo prior to infall. It shines light on a tantalising question – if the Milky Way ($_{\text{200c}}\sim 1\times 10^{12} M, M$$_{\ast}\sim 5\times 10^{10} M$; Bland-Hawthorn & Gerhard, 2016) has over 100 stream candidates with 8 streams spectroscopically confirmed to originate from satellite galaxies, where are the streams disrupting around the LMC ($_{\text{200c}}\sim 2\times 10^{11} M, M$$_{\ast}\sim 3\times 10^{9} M$; Vasiliev, 2023)? Future deep photometry from the Rubin Observatory and Roman Space Telescope could provide exciting answers in the coming decade, potentially adding a new layer to the tangled web of disrupting satellites that comprise the Galactic halo.